OpenStudy (cathyangs):

What is the calculator syntax for logarithms? ex: imputing log base 4 of 16 What is the calculator syntax for logarithms? ex: imputing log base 4 of 16 @Mathematics

OpenStudy (anonymous):

if your calculator has a "log" button, just do:\[\frac{\log 16}{\log 4}\]that will calculate:\[\log_4 16\] for you.

OpenStudy (anonymous):

which should be 2. because:\[\log_4 16 = x \iff 4^x = 16 \iff 4^x=4^2\iff x=2\]

OpenStudy (cathyangs):

ah...thanks. I knew that...and was hoping there was a quicker way. I have a TI-84, and I have to go to catalogue to find log( thanks anyways.

OpenStudy (anonymous):

or you could do:\[\log_4 16=\log_4 4^2 = 2\log_4 4 = 2\cdot 1 = 2\]

OpenStudy (cathyangs):

TI84s dont have that syntax capacity...thank you though.

OpenStudy (anonymous):

The method describeed above is called Change of Bases.